# Control of active field theories at minimal dissipation

### v1

@@ Results
$\dot{\phi}+\nabla\cdot\boldsymbol{J}=0\,,$ (3)
where $\dot{\phi}=\partial_{t}\phi$. We describe the stochastic coupling of $(\phi,n)$ through a linear relation between thermodynamic fluxes $\mathbb{J}=(\boldsymbol{J},\dot{n})$ and forces $\mathbb{F}=(-\nabla(\delta F/\delta\phi),\Delta\mu)$ [38]:
@@
Here, refers to vector transpose, and we have set the Boltzmann constant to unity. Whenever $\mathbb{L}$ depends on $\phi$ and its gradients, the noise $\mathbb{\Lambda}$ is multiplicative. Then, interpreting Eq. (4) within the Stratonovitch convention [59, 60], the spurious drift $-ν\bbnu$ can be explicitly calculated for any given $\mathbb{L}$ [61].
@@
$\mathbb{L}=\begin{pmatrix}\lambda_{\phi}\boldsymbol{1}_{d}&\boldsymbol{C}(\phi,\nabla\phi)\\ \boldsymbol{C}^{\intercal}(\phi,\nabla\phi)&\lambda_{n}\end{pmatrix}\,,$ (6)
@@
We offer a systematic approach to express the heats $(Q_{\rm p},Q_{\rm t})$ as functionals of the protocol $a(t)$ for a generic active field theory [Eqs. (3-6)]. Inspired by previous works [54, 4], we rely on a perturbative expansion around the quasistatic limit ($\Omega\tau\gg 1$, where $\Omega$ is the slowest relaxation frequency of the system [61]), and assume that the protocol is smooth by neglecting any discontinuity in $a(t)$; see dynamical response theory (DRT) in Methods. Within DRT, we obtain an implicit expression of $(Q_{\rm p},Q_{\rm t})$ for an arbitrary protocol $a(t)$ as
@@
where $\langle\cdot\rangle_{\text{s}}$ denotes a stationary average, and $(B_{\rm t},B_{\rm p})$ are boundary terms set by $(a_{0},a_{\tau})$ [61]. The Lagrangian $L$ is the only contribution depending on the full protocol:
@@
Within Lagrangian mechanics, $L$ maps into the Lagrangian of a particle with position $a$ and effective mass $m(a)$ subject to the potential $-\langle P\rangle_{\text{s}}(a)$. We relate $m(a)$ to some response functions (or, alternatively, some correlation functions [4]) measured at fixed $a$ [61].
@@
$F([\phi],a)=\int_{V}\bigg{[}\frac{a(t)}{2}\phi^{2}+\frac{b}{4}\phi^{4}+\frac{\kappa}{2}(\nabla\phi)^{2}\bigg{]}\mathrm{d}x\,,$ (15)
@@
and $J$ is given in Eq. (17). At the mean-field level ($T=0$), the phase-separated profile $\phi(x)$ is known analytically [36], from which we deduce $p(x)$ [61]: it vanishes in the bulk phases, and is negative at the interface [Fig. 2(b) inset]. Therefore, the dissipation is maximal in the homogeneous state, where all the fuel energy is wasted in the thermostat, and is reduced for phase separation, where part of the fuel energy serves to shape the density profile [Fig. 2(b)]. In the homogeneous state, our analytical results [61] show that the finite-temperature value of $\langle P\rangle_{\text{s}}$ is negative, and its absolute value reduces closer to the phase boundary [Fig. 2(a)], while maintaining $W_{\text{act}}\geq 0$.
@@
Figure 3: (a) Protocol heat $Q_{\rm p}$ and (b) total heat $Q_{\rm t}$ as functions of the protocol duration $\Omega\tau$ for the linear protocol $a_{\rm lin}$ [Eq. (21)] at two levels of activity $\Delta\mu/T$. Markers refer to numerical simulations (code available in [68]), solid lines to analytical results, and dashed lines to predictions of DRT [61]. Parameters: $a_{0}=1$, $a_{\tau}=10$, $b=\kappa=1$, $\phi_{0}=0.5$, $\lambda_{\phi}=\lambda_{n}=\gamma=1$, $V/\ell=64$, $F(\phi_{0},a_{0})/T=50V/\ell$.
@@
Using numerical simulations [61], we evaluate the heat $(Q_{\rm p},Q_{\rm t})$ [Eqs. (1-2) and (8-11)] for two levels of activity $\Delta\mu/T=(0.1,0.5)$ [Fig. 3]. In both cases, $Q_{\rm t}$ features a global minimum at an intermediate value of the protocol duration $\tau$. In contrast, $Q_{\rm p}$ has a local minimum for $\Delta\mu/T=0.1$, whereas it increases monotonically for $\Delta\mu/T=0.5$. As previously discussed, these features support the fact that $Q_{\rm t}$ is generally the relevant heat to optimize, since the existence of a minimum is independent of the activity.
@@
To obtain some analytical predictions, we focus on the regime of weak noise where $T$ is much smaller than the local energy scale $(\ell/V)F$, with $\ell$ being the grid spacing, which also corresponds to large system size. Expanding the active field $\phi$ around the mean-field solution $\phi_{0}=\frac{1}{V}\int_{V}\phi\mathrm{d}x$, the dynamics of each mode $\tilde{\phi}_{k}=\frac{1}{V}\int_{V}e^{ikx}(\phi-\phi_{0})\mathrm{d}x$ is decoupled and linear to leading order in $T$. Our corresponding analytical results for $(Q_{\rm p},Q_{\rm t})$ [61] reproduce the curves sampled numerically [Fig. 3]. Alternatively, one can also evaluate $(Q_{\rm p},Q_{\rm t})$ using DRT in the weak-noise regime [61]: these predictions are in good agreement with the numerics at large $\Omega\tau$, yet show a discrepancy at small $\Omega\tau$ [Fig. 3], as expected. Moreover, DRT quantitatively captures the locations of $(\tau_{\rm t},\tau_{\rm p})$ only at a small activity ($\Delta\mu/T=0.1$).
@@
We now analyze how the extremal durations $(\tau_{\rm t},\tau_{\rm p},\tau_{\rm max})$ [Fig. 1] scale with activity $\Delta\mu/T$ and system size $V/\ell$. In particular, our aim is to find the regimes where the DRT predictions are accurate. Based on leading-order expansions [61], we obtain the scalings $(B_{\rm p},m)={\cal O}(V^{2},\Delta\mu^{0})$ and $\left(P_{0},\langle P\rangle_{\text{s}}\right)={\cal O}(V,\Delta\mu^{2})$ at large $V$ and small $\Delta\mu$, from which we deduce the scalings of the minima $(\tau_{\rm p},\tau_{\rm t})={\cal O}(V^{1/2},\Delta\mu^{-1})$ using DRT predictions [Eq. (14)]. Our analytical results for $(\tau_{\rm p},\tau_{\rm t})$ are in good agreement with DRT for small $\Delta\mu/T$ and moderate $V/\ell$ [Figs. 4(a-b)], and the relative errors increase with $(\Delta\mu/T,V/\ell)$ [Figs. 4(c-d)]. Remarkably, the analytical results also suggest the scaling of the maximum $\tau_{\rm max}={\cal O}(V^{2},\Delta\mu^{0})$, from which it follows that the local minimum and maximum of $Q_{\rm p}$ [Fig. 1] always merge at large $(\Delta\mu/T,V/\ell)$ [Figs. 4(a-b)].
@@
We reveal that DRT becomes less accurate as volume $V$ and activity $\Delta\mu$ increase: this result is consistent with the assumption that DRT breaks down when the relaxation time $\Omega^{-1}$, which decreases with $(V,\Delta\mu)$ [61], becomes comparable with $(\tau_{\rm t},\tau_{\rm p},\tau_{\rm max})$. In fact, our DRT scaling predictions are valid for an arbitrary protocol $a(t)$. The scaling $(\tau_{\rm p},\tau_{\rm t})\sim V^{1/2}$ is a consequence of the conservation of particle number: we find a different scaling for active theories of non-conserved fields [61]. In contrast, the scaling $(\tau_{\rm p},\tau_{\rm t})\sim\Delta\mu^{-1}$ is robust: it embodies the competition between internal activity and external perturbation, respectively regulating the small- and large-$\tau$ regimes, inherent to any active system [4].
@@
We continue by examining the DRT predictions for the optimal protocol $a_{\text{op}}(t)$ minimizing $(Q_{\rm p},Q_{\rm t})$ [Eq. (12)] at a given duration $\tau$. This amounts to optimizing the Lagrangian $L$ [Eq. (13)], since the rest of the terms in $Q_{\rm t}$ [Eq. (12)] are independent of the protocol shape $a(t)$. The optimization of $L$ is obtained explicitly by solving the corresponding Euler-Lagrange equation [61]. Since $Q_{\rm p}$ and $Q_{\rm t}$ [Eq. (12)] have the same functional dependence on $a(t)$, we attribute them the same optimal protocol $a_{\rm op}(t)$. Again, we focus here on the CMB dynamics in the weak-noise regime [Eqs. (15-18)] with $a(t)$ evolving within the homogeneous state, see Fig. 5(a).
@@
In practice, Eq. (23) reproduces the value estimated numerically (see [61] for more details) for our specific parameters [Fig. 5(b)]: DRT predictions quantitatively capture the global minimum of $Q_{\rm t}$ whenever their regime of validity contains $\tau_{\rm t}$.
Therefore, we conclude that the crossover between optimal protocols is governed by the competition between the dissipation stemming from either external driving or internal activity, respectively dominant at small and large $\tau$. This crossover is expected to appear generically when controlling homogeneous states in a broad class of active systems, beyond the specific case of CMB [61].
@@ Discussion
For the dynamics of a conserved scalar field that exhibits a nonequilibrium phase-separation, our scaling analysis predicts that $Q_{\rm p}$ does not feature any minimum at large activity $\Delta\mu$ and volume $V$. A similar analysis for non-conserved scalar fields yields a different scaling for $V$ [61], so that $Q_{\rm p}$ may show a local minimum even at large $V$. These results pave the way towards a systematic scaling analysis to inform protocol optimization in a broader class of active field theories. For instance, our approach could be deployed in polar active theories of the flocking transition [40], for which a thermodynamically-consistent description has been recently proposed [49].
@@ Methods
$Q_{\rm t}=\left\langle\int_{0}^{\tau+\tau_{\rm r}}\int_{V}\Delta\mu\,\dot{n}-\frac{\delta F}{\delta\phi}\dot{\phi}\mathrm{d}{\bf r}\mathrm{d}t\right\rangle\,,$ (26)
where we have used the definitions $\mathbb{J}=(\boldsymbol{J},\dot{n})$ and $\mathbb{F}=(-\nabla(\delta F/\delta\phi),\Delta\mu)$. A related expression has been derived for the heat in the absence of perturbation ($\dot{a}=0$) in terms of steady-state averages [38]. For a given protocol $a(t)$, the standard chain rule, valid within the Stratonovitch convention [59], yields
@@
The stochastic integrals in Eq. (29) are interpreted within the Stratonovitch convention. Consequently, we demonstrate in Ref. [61] that $\langle\int P_{\text{spu}}\mathrm{d}t\rangle=0$, so Eq. (10) indeed corresponds to $W_{\rm act}$. In fact, the spurious drift $-ν\bbnu$ ensures thermodynamic consistency of the dynamics [Eq. (4)], but the active work [Eq. (10)] does not explicitly depend on $-ν\bbnu$, as expected for noise-averaged quantities.
@@
Note that the observable-specific quantities in the rhs of Eq. (34), namely $\left\langle O\right\rangle_{\text{s}}$ and $\zeta_{O}^{(m,n)}$, depend only on $a(t)$ and not on its derivatives $(\dot{a},\ddot{a})$. Finally, substituting Eq. (34) into the expression of heat [Eq. (7)] yields the decomposition given in Eq. (12); see [61] for details.

### v2

@@ Results
$\dot{\phi}+\boldsymbol{\nabla}\cdot\boldsymbol{J}=0\,,$ (3)
where $\dot{\phi}=\partial_{t}\phi$. We describe the stochastic coupling of $(\phi,n)$ through a linear relation between thermodynamic fluxes $\mathbb{J}=(\boldsymbol{J},\dot{n})$ and forces $\mathbb{F}=(-\boldsymbol{\nabla}(\delta F/\delta\phi),\Delta\mu)$ [38]:
@@
Here, refers to vector transpose, and we have set the Boltzmann constant to unity. Whenever $\mathbb{L}$ depends on $\phi$ and its gradients, the noise $\mathbb{\Lambda}$ is multiplicative. Then, interpreting Eq. (4) within the Stratonovitch convention [59, 60], the spurious drift $-ν\bbnu$ can be explicitly calculated for any given $\mathbb{L}$ (see Ref. [61], section VI).
@@
$\mathbb{L}=\begin{pmatrix}\lambda_{\phi}\boldsymbol{1}_{d}&\boldsymbol{C}(\phi,\boldsymbol{\nabla}\phi)\\ \boldsymbol{C}^{\intercal}(\phi,\boldsymbol{\nabla}\phi)&\lambda_{n}\end{pmatrix}\,,$ (6)
@@
We offer a systematic approach to express the heats $(Q_{\rm p},Q_{\rm t})$ as functionals of the protocol $a(t)$ for a generic active field theory [Eqs. (3-6)]. Inspired by previous works [54, 4], we rely on a perturbative expansion around the quasistatic limit ($\Omega\tau\gg 1$, where $\Omega$ is the slowest relaxation frequency of the system (see Ref. [61], section IV.B), and assume that the protocol is smooth by neglecting any discontinuity in $a(t)$; see dynamical response theory (DRT) in Methods. Within DRT, we obtain an implicit expression of $(Q_{\rm p},Q_{\rm t})$ for an arbitrary protocol $a(t)$ as
@@
where $\langle\cdot\rangle_{\text{s}}$ denotes a stationary average, and $(B_{\rm t},B_{\rm p})$ are boundary terms set by $(a_{0},a_{\tau})$ (see Ref. [61], section I). The Lagrangian $L$ is the only contribution depending on the full protocol:
@@
Within Lagrangian mechanics, $L$ maps into the Lagrangian of a particle with position $a$ and effective mass $m(a)$ subject to the potential $-\langle P\rangle_{\text{s}}(a)$. We relate $m(a)$ to some response functions (or, alternatively, some correlation functions [4]) measured at fixed $a$ (see Ref. [61], section I).
@@
$F([\phi],a)=\int_{V}\bigg{[}\frac{a(t)}{2}\phi^{2}+\frac{b}{4}\phi^{4}+\frac{\kappa}{2}(\boldsymbol{\nabla}\phi)^{2}\bigg{]}\mathrm{d}x\,,$ (15)
@@
and $J$ is given in Eq. (17). At the mean-field level ($T=0$), the phase-separated profile $\phi(x)$ is known analytically [36], from which we deduce $p(x)$ (see Ref. [61], section IV.A): it vanishes in the bulk phases, and is negative at the interface [Fig. 2(b) inset]. Therefore, the dissipation is maximal in the homogeneous state, where all the fuel energy is wasted in the thermostat, and is reduced for phase separation, where part of the fuel energy serves to shape the density profile [Fig. 2(b)]. In the homogeneous state, our analytical results (see Ref. [61], section IV.B) show that the finite-temperature value of $\langle P\rangle_{\text{s}}$ is negative, and its absolute value reduces closer to the phase boundary [Fig. 2(a)], while maintaining $W_{\text{act}}\geq 0$.
@@
Figure 3: (a) Protocol heat $Q_{\rm p}$ and (b) total heat $Q_{\rm t}$ as functions of the protocol duration $\Omega\tau$ for the linear protocol $a_{\rm lin}$ [Eq. (21)] at two levels of activity $\Delta\mu/T$. Markers refer to numerical simulations (code available in [68]), solid lines to analytical results, and dashed lines to predictions of DRT [Eqs. (12)]. Parameters: $a_{0}=1$, $a_{\tau}=10$, $b=\kappa=1$, $\phi_{0}=0.5$, $\lambda_{\phi}=\lambda_{n}=\gamma=1$, $V/\ell=64$, $F(\phi_{0},a_{0})/T=50V/\ell$.
@@
Using numerical simulations (see Ref. [61], section IV.G), we evaluate the heat $(Q_{\rm p},Q_{\rm t})$ [Eqs. (1-2) and (8-11)] for two levels of activity $\Delta\mu/T=(0.1,0.5)$ [Fig. 3]. In both cases, $Q_{\rm t}$ features a global minimum at an intermediate value of the protocol duration $\tau$. In contrast, $Q_{\rm p}$ has a local minimum for $\Delta\mu/T=0.1$, whereas it increases monotonically for $\Delta\mu/T=0.5$. As previously discussed, these features support the fact that $Q_{\rm t}$ is generally the relevant heat to optimize, since the existence of a minimum is independent of the activity.
@@
To obtain some analytical predictions, we focus on the regime of weak noise where $T$ is much smaller than the local energy scale $(\ell/V)F$, with $\ell$ being the grid spacing, which also corresponds to large system size. Expanding the active field $\phi$ around the mean-field solution $\phi_{0}=\frac{1}{V}\int_{V}\phi\mathrm{d}x$, the dynamics of each mode $\tilde{\phi}_{k}=\frac{1}{V}\int_{V}e^{ikx}(\phi-\phi_{0})\mathrm{d}x$ is decoupled and linear to leading order in $T$. Our corresponding analytical results for $(Q_{\rm p},Q_{\rm t})$ (see Ref. [61], section III.B) reproduce the curves sampled numerically [Fig. 3]. Alternatively, one can also evaluate $(Q_{\rm p},Q_{\rm t})$ using DRT in the weak-noise regime (see Ref. [61], section III): these predictions are in good agreement with the numerics at large $\Omega\tau$, yet show a discrepancy at small $\Omega\tau$ [Fig. 3], as expected. Moreover, DRT quantitatively captures the locations of $(\tau_{\rm t},\tau_{\rm p})$ only at a small activity ($\Delta\mu/T=0.1$).
@@
We now analyze how the extremal durations $(\tau_{\rm t},\tau_{\rm p},\tau_{\rm max})$ [Fig. 1] scale with activity $\Delta\mu/T$ and system size $V/\ell$. In particular, our aim is to find the regimes where the DRT predictions are accurate. Based on leading-order expansions (see Ref. [61], section IV.D), we obtain the scalings $(B_{\rm p},m)={\cal O}(V^{2},\Delta\mu^{0})$ and $\left(P_{0},\langle P\rangle_{\text{s}}\right)={\cal O}(V,\Delta\mu^{2})$ at large $V$ and small $\Delta\mu$, from which we deduce the scalings of the minima $(\tau_{\rm p},\tau_{\rm t})={\cal O}(V^{1/2},\Delta\mu^{-1})$ using DRT predictions [Eq. (14)]. Our analytical results for $(\tau_{\rm p},\tau_{\rm t})$ are in good agreement with DRT for small $\Delta\mu/T$ and moderate $V/\ell$ [Figs. 4(a-b)], and the relative errors increase with $(\Delta\mu/T,V/\ell)$ [Figs. 4(c-d)]. Remarkably, the analytical results also suggest the scaling of the maximum $\tau_{\rm max}={\cal O}(V^{2},\Delta\mu^{0})$, from which it follows that the local minimum and maximum of $Q_{\rm p}$ [Fig. 1] always merge at large $(\Delta\mu/T,V/\ell)$ [Figs. 4(a-b)].
@@
We reveal that DRT becomes less accurate as volume $V$ and activity $\Delta\mu$ increase: this result is consistent with the assumption that DRT breaks down when the relaxation time $\Omega^{-1}$, which decreases with $(V,\Delta\mu)$ (see Ref. [61], section IV.B), becomes comparable with $(\tau_{\rm t},\tau_{\rm p},\tau_{\rm max})$. In fact, our DRT scaling predictions are valid for an arbitrary protocol $a(t)$. The scaling $(\tau_{\rm p},\tau_{\rm t})\sim V^{1/2}$ is a consequence of the conservation of particle number: we find a different scaling for active theories of non-conserved fields (see Ref. [61], section V). In contrast, the scaling $(\tau_{\rm p},\tau_{\rm t})\sim\Delta\mu^{-1}$ is robust: it embodies the competition between internal activity and external perturbation, respectively regulating the small- and large-$\tau$ regimes, inherent to any active system [4].
@@
We continue by examining the DRT predictions for the optimal protocol $a_{\text{op}}(t)$ minimizing $(Q_{\rm p},Q_{\rm t})$ [Eq. (12)] at a given duration $\tau$. This amounts to optimizing the Lagrangian $L$ [Eq. (13)], since the rest of the terms in $Q_{\rm t}$ [Eq. (12)] are independent of the protocol shape $a(t)$. The optimization of $L$ is obtained explicitly by solving the corresponding Euler-Lagrange equation (see Ref. [61], section IV.E). Since $Q_{\rm p}$ and $Q_{\rm t}$ [Eq. (12)] have the same functional dependence on $a(t)$, we attribute them the same optimal protocol $a_{\rm op}(t)$. Again, we focus here on the CMB dynamics in the weak-noise regime [Eqs. (15-18)] with $a(t)$ evolving within the homogeneous state, see Fig. 5(a).
@@
In practice, Eq. (23) reproduces the value estimated numerically (see Ref. [61], section IV.F for more details) for our specific parameters [Fig. 5(b)]: DRT predictions quantitatively capture the global minimum of $Q_{\rm t}$ whenever their regime of validity contains $\tau_{\rm t}$.
Therefore, we conclude that the crossover between optimal protocols is governed by the competition between the dissipation stemming from either external driving or internal activity, respectively dominant at small and large $\tau$. This crossover is expected to appear generically when controlling homogeneous states in a broad class of active systems, beyond the specific case of CMB (see Ref. [61], section IV.E).
@@ Discussion
For the dynamics of a conserved scalar field that exhibits a nonequilibrium phase-separation, our scaling analysis predicts that $Q_{\rm p}$ does not feature any minimum at large activity $\Delta\mu$ and volume $V$. A similar analysis for non-conserved scalar fields yields a different scaling for $V$ (see Ref. [61], section V), so that $Q_{\rm p}$ may show a local minimum even at large $V$. These results pave the way towards a systematic scaling analysis to inform protocol optimization in a broader class of active field theories. For instance, our approach could be deployed in polar active theories of the flocking transition [40], for which a thermodynamically-consistent description has been recently proposed [49].
@@ Methods
$Q_{\rm t}=\left\langle\int_{0}^{\tau+\tau_{\rm r}}\int_{V}\bigg{(}\Delta\mu\,\dot{n}-\frac{\delta F}{\delta\phi}\dot{\phi}\bigg{)}\mathrm{d}{\bf r}\mathrm{d}t\right\rangle\,,$ (26)
where we have used the definitions $\mathbb{J}=(\boldsymbol{J},\dot{n})$ and $\mathbb{F}=(-\boldsymbol{\nabla}(\delta F/\delta\phi),\Delta\mu)$. A related expression has been derived for the heat in the absence of perturbation ($\dot{a}=0$) in terms of steady-state averages [38]. For a given protocol $a(t)$, the standard chain rule, valid within the Stratonovitch convention [59], yields
@@
The stochastic integrals in Eq. (29) are interpreted within the Stratonovitch convention. Consequently, we demonstrate in Ref. [61], section VI that $\langle\int P_{\text{spu}}\mathrm{d}t\rangle=0$, so Eq. (10) indeed corresponds to $W_{\rm act}$. In fact, the spurious drift $-ν\bbnu$ ensures thermodynamic consistency of the dynamics [Eq. (4)], but the active work [Eq. (10)] does not explicitly depend on $-ν\bbnu$, as expected for noise-averaged quantities.
@@
Note that the observable-specific quantities in the rhs of Eq. (34), namely $\left\langle O\right\rangle_{\text{s}}$ and $\zeta_{O}^{(m,n)}$, depend only on $a(t)$ and not on its derivatives $(\dot{a},\ddot{a})$. Finally, substituting Eq. (34) into the expression of heat [Eq. (7)] yields the decomposition given in Eq. (12); see Ref. [61], section I for details.